OpenStudy (anonymous):

\[f(x)=\frac 1 x , a=-3\]

OpenStudy (anonymous):

Are you sure You will write it today??

OpenStudy (anonymous):

 I made a table so now I just gotta write the Taylor series \[f(x)=\sum_{n=0}^{\infty}\frac{f^n a}{n!}(x-a)^n\] \[f(-3)+\frac{f'(3)}{1!}(x-3)+\frac{f''(3)}{2!}(x-3)^2+.....\] I don't know what the heck I'm doing!!!!! Yeah!!!! It's 4:36am where I'm from!!!! Yeah!!! ....ok I just lost it LOL